Air Travel
An airplane flew for 4 hr with a 20-mph tailwind. the return flight against the same wind took 5 hr. Find the speed of the plane in still air
d + r (t)

distance rate time
with wind r + 20
d


Answer by edjones(8007)

d=rt
rt=rt because the distance is the same coming and going.
Let r be the speed in still air.
4(r+20)=5(r-20)
4r+80=5r-100
-r=-180
r=180 mph
.
Ed